OpenStudy (anonymous):

1/4(3l+27)=3/4(3l+15)

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

There is no variable to solve for?

OpenStudy (anonymous):

its L

OpenStudy (anonymous):

We begin by distributing the 4 in both the denominators This yields: 1/12l+108=3/12l+60. Next we multiply both sides by 12l+60. This gives us: 12l+60/12l+108=3. Next we multiply both sides by 12l+108. This results in; 12l+60=3(12l+108). After distributing the 3 we get, 12l+60=36l+324. We subtract 60 from both sides. We get 12l=36l+264. Next subtract 36l from both sides, this gives us -24l=264. Divide both sides by -24 and we get l=-11

OpenStudy (anonymous):

that helped alot thanx